# Performance Improvement of Deorbitalized Exchange-Correlation Functionals

### v1

@@ IV.2 Timing results
Table 4 presents timing statistics for the 223 molecule G3X/99 test set. Table 5 displays the timing data for the six-molecule subset, denoted AE6, of the G3 test set. Each table shows the total time for the run of a test set, and the average time taken per system. In addition, each table shows the average number of SCF cycles needed to converge to self-consistency, (defined as the total number of SCF cycles divided by the number of systems) and the average time per SCF cycle, taken as the average time divided by the average number of cycles. In Table 4, spreads in these last two measures are shown also. Those are defined, as before, as the difference between the maximum and minimum values of a quantity observed across a data set.
@@
Table 4 shows that both SRPP and SRPP2 are substantially faster than RPP for the molecules. Similar to the PCopt  deorbitalization, neither of them delivers a speed advantage over the parent, orbital-dependent functional. The effects of test-set sampling are shown in Table 5. The AE6 timings give a much more favorable comparison of SRPP2 over RPP but no meaningful gain for SRPP2 versus either PCopt  or the parent r2SCAN functional.
@@
Methodological effects are shown in Table 6. It gives the timings for the G3 test with the VASP 5.4.4 calculations done as isolated systems in a large orthorhombic box. Distinct from the solid calculations, the default cutoff energy was set to $600$ eV and $\Gamma$ point sampling was used. Unlike the G3 calculations in NWChem, the VASP tests exhibit a large degradation in performance by both RPP and SRPP deorbitalizations with respect to the parent functional. Interestingly, the time per SCF cycle for the SRPP deorbitalizer is 1.29 s/cycle in VASP, while in NWChem it is 1.28 s/cycle. However, for the parent functional r2SCAN, VASP takes 2.87 s/cycle compared to 1.34 s/cycle in NWChem. Unfortunately, the number of SCF cycles is much higher for the deorbitalized forms, roughly a factor of 2.5 for SRPP and a factor of 3 for RPP. In NWChem the average number of SCF cycles is roughly the same for all three. The speed advantage per cycle of the deorbitalized forms is lost thereby in the quasi-molecular case. This brings to light the drastic effect basis set methodology (plane-wave PAW versus gaussian) can have on the performance of the deorbitalized forms.
@@
Table 6: As in Table 4 for the 223 molecules, but with the calculations done in VASP 5.4.4. for the deorbitalized functionals RPP and SRPP as compared to the parent functional.
@@
For solids, Table 7 presents a clearly different story. Both SRPP and SRPP2 outperform the parent, orbital-dependent functional on total time. SRPP2 in particular needs only about 75% of the time of r2SCAN. Notably, SRPP2 also requires only about 72% more time than PBE. This overall gain compared to the PCopt  deorbitalization comes from a drastic reduction in the number of SCF cycles needed by SRPP and SRPP2, about half of the number required by PCopt. Importantly, the times per SCF cycle for SRPP and SRPP2 are essentially the same as for RPP, less than half that for r2SCAN, and actually faster than even for PBE. The combined result is an overall speedup.
@@
Fig. 5(b) shows the number of SCF cycles needed for r2SCAN and the three deorbitalized versions considered in the preceding plot. One sees clearly how the SCF process for original r2SCAN converges in consistently fewer cycles than for any of the three deorbitalized variants. The RPP-type deorbitalizers perform quite reasonably but with a number of outliers. In contrast PCopt  has a rather larger number of outliers which cancel its edge in the time-per-cycles metric. Note also the scatter plot of SCF cycles needed versus the r2SCAN requirement in Fig. 6. The smoothed SRPP and SRPP2 do provide consistent improvement upon the performance of RPP, with a few exceptions among the semiconductors and simple metals.
@@
IV.3 Detailed analysis of computational parameter dependence
@@
Table 8: Parameter and SCF algorithm dependence of timing for computing the set of $5$ solids, for r2SCAN itself, and r2SCAN-L deorbitalized with RPP or SRPP. $E_{\mathrm{cut}}$ is the plane-wave cut-off energy; $E_{\mathrm{diff}}$ is the energy tolerance used to end the self-consistency cycle. The algorithms to solve the KS or gKS equation, using the keyword terminology for the ALGO input parameter in VASP, are: (All) preconditioned conjugate gradient, (Normal) blocked Davidson iteration, (Fast) hybrid of block-Davidson and RMM-DIIS (Residual Minimization Method with Direct Inversion in the Iterative Subspace), and (Very Fast) RMM-DIIS.
@@
In contrast, the advantage in terms of the number of cycles needed to reach self-consistency for the gKS procedure used with r2SCAN over the deorbitalized forms and their pure KS potential does vary quite a bit depending on the tightness of convergence criteria. gKS loses its advantage relative to deorbitalized KS as convergence is loosened and SRPP loses its advantage relative to RPP. Ultimately, for Case 3, the sloppiest one for preconditioned conjugate gradients and Case 6, the sloppiest overall, there is basically no difference in performance among the three strategies in this regard. Therefore the total time speed-up of deorbitalization is quite dramatic, almost a factor of 2. (It may seem peculiar that all functionals in Case 6 take substantially more cycles to achieve self-consistency than Case 3, with both larger plane-wave cutoff and stricter convergence tolerance. This is a result of the instability of the RMS-DIIS, which makes sacrifices in stability to achieve faster overall times Kresse .
There is a very plausible, simple explanation for such a finding. Reducing the cut-off energy amounts to introducing a low-pass filter that eliminates a good deal of noise in the potential. Though the RPP and SRPP potentials have very different noise characteristics, it is reasonable to suppose that those differences are substantially removed by smoothing with a low energy cutoff. Similarly, a less demanding tolerance means greater acceptance of effects of noise in the potential, reducing their consequences. We hypothesize that the smooth atomic-like basis used in molecular calculations is coarse relative to individual plane waves, therefore causing a similar noise reduction in the potential and yielding the “egalitarian" timing results seen for these calculations.
IV.4 Molecular dynamics of Aluminum
@@
Solid Al is a particularly advantageous case for deorbitalization. Consistent with the discussion above about the 55 solid test set (recall Table 7), all three deorbitalizations (PCopt, RPP, SRPP) are substantially faster in total time than the parent functional. The spreads are even better for RPP and PCopt, and as good as the parent for SRPP. Also as expected, RPP and SRPP outperform PCopt substantially with regard to the required number of SCF cycles.
@@
The simulation system consisted of 108 atoms, at bulk density $\rho=2.34~\mathrm{g/cm}^{3}$ in a $12.7239\times 12.7239\times 12.7239$ Å3 cell. The $\Gamma$-point was used to sample the Brillouin zone. In VASP 5.4.4, the system was treated with a three-electron pseudopotential ($3s^{2}$ $3p^{1}$ valence, 10 electrons in the core; denoted PAW_PBE Al_GW). Non-spherical contributions within the PAW spheres were included self-consistently. Minimization used the RMM-DISS algorithm with energy cutoff set to 500 eV, as in Case 6 in Table 8. The self-consistent energy convergence tolerance $E_{\mathrm{diff}}$ was $1e^{-4}$ eV. Approximate Fermi-type thermal smearing with a width of 0.0881555 eV was applied, and the initial temperature set to 1023 K. Molecular dynamics used the Verlet algorithm and a Nosé thermostat. 6501 steps of time-step 0.941 femtoseconds were taken for each simulation. Each calculation utilized 5 nodes on the University of Florida HiperGator cluster (Gen. 3), with 64 cores per node (total of 320 cores) and 4 GB of memory per core (1280 GB total).
@@
One way to characterize this problematic cycle count is that convergence to self-consistency from the fully converged density of the previous MD step is not that much quicker than convergence from an LCAO density (12 cycles average per MD step for SRPP, in contrast to 14 cycles average over the test set and 18 to 21 for the equation of state calculation for Al). The behavior also is in contrast with the dramatic reduction of PBE and r2SCAN requirements to less than 3 cycles average per step. This may be an unpleasant consequence of the inclusion of the density Laplacian in the X functional and associated sensitivity to seemingly small changes in the density. A peculiarity is that a tolerance of $E_{\mathrm{diff}}=2.75\times 10^{-5}$ needs an average of 14 cycles to converge in the case of the 55 element test set (see Table 8) while, on the other hand, use of $E_{\mathrm{diff}}=1\times 10^{-4}$ in the MD needs 18 cycles. The difficulty is that though RMS-DIIS is extremely fast it is less stable than the other solvers.
@@ V Discussion and Concluding remarks
We have constructed a measure, $I^{deorb}$, to control instability sources in the XC potential of deorbitalized metaGGAs, in terms of a metric derived from the Pauli potential (the functional derivative of the deorbitalized kinetic energy density with respect to the density). This measure clearly diagnoses the relative instability or reliability of performance of standard deorbitalizers and helps motivate a modified form of RPP, the SRPP and SRPP2.
@@
In consequence, SRPP and SRPP2 produce noticeably smoother kinetic and exchange potentials than other deorbitalizers. Both modified forms preserve or modestly enhance the predictive performance of RPP in the deorbitalization of the r2SCAN functional on a standard test set for the structural properties of solids. But for the smoothed functionals, this constraint compliance for slowly varying densities comes with a much smaller penalty in predictive performance for highly inhomogeneous systems than one might expect from the behavior of the RPP. The smoothed functionals improve upon RPP markedly for the G3 molecular test set, though the deorbitalized form still is not competitive with r2SCAN-L(PCopt) for heats of formation. The SRPP2 variant, though less effective than SRPP on the molecular cases, is equal or better on the solids and delivers better timing performance. It is plausible that this gain is a consequence of the SRPP2 X potential being smoother than the SRPP one.
The implications for timing performance are complicated and unfortunately somewhat obscure. For molecules, both the time per SCF cycle and the number of cycles needed to reach convergence are essentially the same for both the parent and deorbitalized functionals.
@@
For fixed (static) solid geometries, the story is different. Unlike earlier deorbitalized functionals, these particular deorbitalized metaGGAs significantly outperform the explicitly orbital dependent, metaGGA parent (used in gKS form) so far as time per SCF cycle is concerned. While the deorbitalized forms (r2SCAN-L with L either SRPP or SRPP2) need more cycles to reach self-consistent convergence, the increment is not so large as to offset the much shorter cycle time. As a result, SRPP and SRPP2 are much faster than r2SCAN for static lattice solid calculations, as seen in Tables 7, 8 and 9.
@@
That advantage does not propagate into the one AIMD example we have tried. It appears that the density Laplacian dependence in the deorbitalized functionals introduces instabilities that greatly increase the number of SCF cycles needed in the AIMD context. Our speculation has been that since the XC potential is sensitive to small perturbations in the density because of the Laplacian, potentials for successive AIMD steps may occasionally be different enough to lead to unusual variation in the density, and thus slower SCF convergence. However, the limited testing we have done with reduced AIMD step-time is not consistent with this argument. Similarly, one might think that there could be a clue as to the cause of this behavior in the average SCF cycle time versus its spread. But Table 7 shows that SRPP, SRPP2, and RPP itself all have similar ratios of spread of time per cycle to average time per cycle. In fact, the ratio for SRPP (2.35) is slightly below that for RPP (2.51).
In the case of plane-wave pseudopotential calculations of solids, we see two outcomes for equation-of-state calculations:
  1. 1.
    For high accuracy calculations, the performance of deorbitalized metaGGA XC functionals relative to the parent is markedly dependent on the smoothness of the the XC potential, with smoothed functionals (SRPP and SRPP2) significantly outperforming not only the earlier deorbitalizers (e.g. PCopt), but the parent orbital-dependent functional itself. In that regard, this work has achieved one of the long-standing goals of the deorbitalization endeavor.
  2. 2.
    Reducing accuracy requirements via either lowered energy cut-off for orbital representations or the energy tolerance for achieving self-consistency, causes all the metaGGA forms (orbital or deorbitalized) to reach rough parity of the number of cycles needed to achieving self-consistency. The orbital-dependent gKS calculations continue to be inferior compared to pure KS from the deorbitalized functionals on a per-cycle basis. But the relative advantage of SRPP and SRPP2 versus the other deorbitalizers is largely lost because lowering the cutoff effectively smooths all potentials to the same degree, while reducing the SCF tolerance reduces the sensitivity of the calculation to instabilities from the potential, so the wall-clock times for all tend to converge.
@@
Because Gaussian basis functions are themselves smooth and are equivalent to fixed combinations of many plane waves, molecules calculated with them behave, oddly enough, like the cruder and faster plane-wave calculations, with little distinction in timing among forms of deorbitalizers. Apparently, the nature of limiting the basis has a counterpart smoothing, hence stabilizing effect as does reducing the energy cutoff in a plane-wave code.
@@
AIMD calculations provide a third, somewhat discomfiting pattern of behavior. As for the case of lower accuracy calculations for the equation of state, deorbitalized functions outperform the orbital-dependent parent on a per-cycle basis, while smoothed and unsmoothed deorbitalizers have similar SCF convergence rates. But PBE and r2SCAN perform the task specific to MD, namely finding orbitals for a new set of nuclear positions starting from the density for a modestly different set of positions, in two or three SCF cycles. In contrast, the deorbitalized forms used here take almost the same number of cycles as they do when starting from, say, an LCAO density. An attempt to start at each position with a PBE density did not help.
One ansatz for AIMD that remains to be investigated would be to take the density from some numerically stable method (PBE to keep it simple), and evaluate the total energy and forces with a single non-self-consistent calculation with the metaGGA-L functional of interest. For the comparatively coarse convergence levels required, that should give nearly the same forces (and energies) as from the conventional self-consistent metaGGA-L, and better than those from self-consistent PBE, at a fraction of the timing cost of self-consistent metaGGA-L.) Doing so would combine best-time performance vis-a-vis finding the density at any step with something that is fast (for a single SCF step) and accurate to calculate forces.
@@
As another line of pursuit, it may be advantageous to take the measure of noise defined in this paper and use it as an aid to construct deorbitalizers with yet smoother potentials, either by optimizing the switching function between iso-orbital and slowly-varying limits or strategically modifying “appropriate norms" in the deorbitalizer or parent functional. This is the focus of ongoing research.
@@
The following file is available free of charge.
  • SRPP_Timing.SuppInfo: In this file, we provide a detailed, system-by-system tabulation of the numerical results of the test calculations against standard molecular and solid test sets. We also provide system-by system SCF timing and SCF cycle count tabulations, corresponding to the plot of SCF cycle count for each of the first 100 AIMD steps in Fig. 8 above.

### v2

@@ IV.3 Timing results
Table 4 presents timing statistics for the 223 molecule G3X/99 test set. Table 5 displays the timing data for the six-molecule subset, denoted AE6, of the G3 test set. Each table shows the total time for the run of a test set, and the average time taken per system. In addition, each table shows the average number of SCF cycles needed to converge to self-consistency, (defined as the total number of SCF cycles divided by the number of systems) and the average time per SCF cycle, taken as the average time divided by the average number of cycles. In Table 4, spreads in these last two measures are shown also. As before, spreads are defined as the difference between the maximum and minimum values of a quantity observed across a data set.
@@
Table 4 shows that both r2SCAN-L(SRPP) and (SRPP2) are substantially faster than r2SCAN-L(RPP) for the molecules. Similar to the PCopt  deorbitalization, neither of them delivers a speed advantage over the parent, orbital-dependent functional. The effects of test-set sampling are shown in Table 5. The AE6 timings give a much more favorable comparison of SRPP2 over RPP but no meaningful gain for SRPP2 versus either PCopt  or the parent r2SCAN functional.
@@
Methodological effects are shown in Table 6. It gives the timings for the G3 test with the VASP 5.4.4 calculations done as isolated systems in a large orthorhombic box. Distinct from the solid calculations, the default cutoff energy was set to $600$ eV and $\Gamma$ point sampling was used. Unlike the G3 calculations in NWChem, the VASP tests exhibit a large degradation in performance by both RPP and SRPP deorbitalizations with respect to the parent functional. Interestingly, the time per SCF cycle for the SRPP deorbitalizer is 1.29 s/cycle in VASP, while in NWChem it is 1.28 s/cycle. However, for the parent functional r2SCAN, VASP takes 2.87 s/cycle compared to 1.34 s/cycle in NWChem. Unfortunately, the number of SCF cycles is much higher for the deorbitalized forms, roughly a factor of 2.5 for SRPP and a factor of 3 for RPP. In NWChem the average number of SCF cycles is roughly the same for all three. The speed advantage per cycle of the deorbitalized forms is lost thereby in the quasi-molecular case. This brings to light the drastic effect that basis set methodology (plane-wave PAW versus gaussian) can have on the performance of the deorbitalized forms.
@@
Table 6: As in Table 4 for the 223 molecules, but with the calculations done in VASP 5.4.4 for the functionals deorbitalized with RPP and SRPP as compared to the parent functional.
@@
For solids, Table 7 presents a clearly different story. r2SCAN deorbitalized with SRPP or SRPP2 outperforms the parent, orbital-dependent functional on total time. SRPP2 in particular needs only about 75% of the time of r2SCAN. Notably, SRPP2 also requires only about 72% more time than PBE. This overall gain compared to the PCopt  deorbitalization comes from a drastic reduction in the number of SCF cycles needed by SRPP and SRPP2, about half of the number required by PCopt. Importantly, the times per SCF cycle for SRPP and SRPP2 are essentially the same as for RPP, less than half that for r2SCAN, and actually faster than even for PBE. The combined result is an overall speedup.
@@
Fig. 5(b) shows the number of SCF cycles needed for r2SCAN and the three deorbitalized versions considered in the preceding plot. One sees clearly how the SCF process for original r2SCAN converges in consistently fewer cycles than for any of the three deorbitalized variants. The RPP-type deorbitalizers perform quite reasonably but with a number of outliers. In contrast PCopt  has a rather larger number of outliers which cancel its edge in the time-per-cycles metric. Note also the scatter plot of SCF cycles needed versus the r2SCAN requirement in Fig. 6. The smoothed r2SCAN-L(SRPP) and (SRPP2) do provide consistent improvement upon the performance of r2SCAN-L(RPP), with a few exceptions among the semiconductors and simple metals.
@@
IV.4 Detailed analysis of computational parameter dependence
@@
Table 8: Parameter and SCF algorithm dependence of timing for computing the set of 55 solids, for r2SCAN itself, and r2SCAN-L deorbitalized with RPP or SRPP. $E_{\mathrm{cut}}$ is the plane-wave cut-off energy; $E_{\mathrm{diff}}$ is the energy tolerance used to end the self-consistency cycle. The algorithms to solve the KS or gKS equation, using the keyword terminology for the ALGO input parameter in VASP, are: (All) preconditioned conjugate gradient, (Normal) blocked Davidson iteration, (Fast) hybrid of block-Davidson and RMM-DIIS (Residual Minimization Method with Direct Inversion in the Iterative Subspace), and (Very Fast) RMM-DIIS.
@@
In contrast, the advantage in terms of the number of cycles needed to reach self-consistency for the gKS procedure used with r2SCAN over the deorbitalized forms and their pure KS potential does vary quite a bit depending on the tightness of convergence criteria. gKS loses its advantage relative to deorbitalized KS as convergence is loosened and SRPP loses its advantage relative to RPP. Ultimately, for Case 3, the sloppiest one for preconditioned conjugate gradients and Case 6, the sloppiest overall, there is basically no difference in performance among the three strategies in this regard. Therefore the total time speed-up of deorbitalization is quite dramatic, almost a factor of 2. (It may seem peculiar that all functionals in Case 6 take substantially more cycles to achieve self-consistency than Case 3, with both larger plane-wave cutoff and stricter convergence tolerance. This is a result of the instability of the RMM-DIIS, which makes sacrifices in stability to achieve faster overall times [58].
There is a plausible, simple explanation for such a finding. Reducing the cut-off energy amounts to introducing a low-pass filter that eliminates a good deal of noise in the potential. Though the r2SCAN-L(RPP) and (SRPP) potentials have very different noise characteristics, it is reasonable to suppose that those differences are substantially removed by smoothing with a low energy cutoff. Similarly, a less demanding tolerance means greater acceptance of effects of noise in the potential, reducing their consequences. We hypothesize that the smooth atomic-like basis used in molecular calculations is coarse relative to individual plane waves, therefore causing a similar noise reduction in the potential and yielding the “egalitarian" timing results seen for these calculations.
IV.5 Molecular dynamics of Aluminum
@@
Solid Al is a particularly advantageous case for deorbitalization. Consistent with the discussion above about the 55 solid test set (recall Table 7), all three deorbitalizations (PCopt, RPP, SRPP) are substantially faster in total time than the parent functional. The spreads are even better for r2SCAN-L(RPP) and (PCopt), and as good as the parent for (SRPP). Also as expected, the RPP- and SRPP-deorbitalized functionals outperform PCopt substantially with regard to the required number of SCF cycles.
@@
The simulation system consisted of 108 atoms, at bulk density $\rho=2.34~\mathrm{g/cm}^{3}$ in a $12.7239\times 12.7239\times 12.7239$ Å3 cell. The $\Gamma$-point was used to sample the Brillouin zone. In VASP 5.4.4, the system was treated with a three-electron PAW effective potential ($3s^{2}$ $3p^{1}$ valence, 10 electrons in the core; denoted PAW_PBE Al_GW). Non-spherical contributions within the PAW spheres were included self-consistently. Minimization used the RMM-DISS algorithm with energy cutoff set to 500 eV, as in Case 6 in Table 8. The self-consistent energy convergence tolerance $E_{\mathrm{diff}}$ was $1\times 10^{-4}$ eV. Approximate Fermi-type thermal smearing with a width of 0.0881555 eV was applied, and the initial temperature set to 1023 K. The molecular dynamics was done in the canonical (NVT) ensemble using the Verlet algorithm and a Nosé thermostat. 6501 steps with time-step 0.941 femtoseconds were taken for each simulation. Each calculation utilized 5 nodes on the University of Florida HiperGator cluster (Gen. 3), with 64 cores per node (total of 320 cores) and 4 GB of memory per core (1280 GB total).
@@
One way to characterize this problematic cycle count is that convergence to self-consistency from the fully converged density of the previous MD step is not that much quicker than convergence from an LCAO density (12 cycles average per MD step for SRPP, in contrast to 14 cycles average over the test set and 18 to 21 for the equation of state calculation for Al). The behavior also is in contrast with the dramatic reduction of PBE and r2SCAN requirements to less than 3 cycles average per step. This may be an unpleasant consequence of the inclusion of the density Laplacian in the X functional and associated sensitivity to seemingly small changes in the density. A peculiarity is that a tolerance of $E_{\mathrm{diff}}=2.75\times 10^{-5}$ needs an average of 14 cycles to converge in the case of the 55 element test set (see Table 8) while, on the other hand, use of $E_{\mathrm{diff}}=1\times 10^{-4}$ in the MD needs 18 cycles. The difficulty is that though RMM-DIIS is extremely fast it is less stable than the other solvers.
@@
We also analyzed the Hellmann-Feynman forces during the course of this simulation. The detailed analysis is provided in Fig. 4 of the Supplemental Information, which shows the maximum, minimum, and median forces in the z direction for the first 50 and 500 MD steps. No outliers are observed, and the median values remain close to zero for all functionals tested. That indicates a symmetric distribution of forces around zero, without systematic bias introduced by the deorbitalized functionals.
@@ V Discussion and Concluding remarks
Deorbitalizing a conventional meta-GGA XC functional is motivated by two closely related goals. First is to regain the interpretability provided by the locality of the KS potential. Second is to achieve major computational speedups as a consequence of that simplification without compromise of predictive accuracy. Thus, the simultaneous focus of this paper is on issues of both functional accuracy and speed. In that regard, Laplacian-dependent deorbitalizers are two-edged swords. While providing useful deorbitalization accuracy, they can introduce highly spiky functions that can be impediments to computational speed.
We have addressed those competing effects by introducing a smoothness measure, $I^{deorb}$, for an XC potential to motivate the use of deorbitalizers (SRPP and SRPP2) designed with smoothness in mind. Both preserve fundamental constraints respected by the recent OFR2 deorbitalization of r2SCAN and do not degrade norm satisfaction significantly. The result is improvement for both accuracy and timing objectives. Our timing study concentrates on approaches (gaussian basis sets, plane-waves and PAW pseudopotentials) used in AIMD and high through-put calculations, and should provide ballpark expectations for any such code. Tuning or adapting these codes for optimal performance for Laplacian-based functionals, especially for AIMD, would be a welcome next step to achieve the full promise of deorbitalization.
@@
We have used the smoothness measure, $I^{deorb}$, to probe for and control instability sources in the XC potential of deorbitalized meta-GGAs. In consequence, SRPP and SRPP2 produce noticeably smoother kinetic and exchange potentials than other deorbitalizers. Notably, for these smoothed variants of RPP, the constraint compliance for slowly varying densities imposed by the RPP comes with a much smaller penalty in predictive performance for highly inhomogeneous systems than one might expect from the behavior of the RPP. The smoothed functionals improve upon RPP markedly for the G3 molecular test set, though the deorbitalized form still is not competitive with r2SCAN-L(PCopt) for heats of formation. The (SRPP2) variant, though less effective than (SRPP) on the molecular cases, is equal or better on the solids and delivers better timing performance. It is plausible that this gain is a consequence of the r2SCAN-L(SRPP2) X potential being smoother than the SRPP-deorbitalized one. It also is plausible that smoothing along the lines we have presented would be helpful in reaching desired performance goals for deorbitalized finite-temperature functionals [59, 60].
@@
The implications for timing performance are complicated and unfortunately somewhat obscure. For molecules, both the time per SCF cycle and the number of cycles needed to reach convergence are essentially the same for both the parent and deorbitalized functionals. For fixed (static) solid geometries, however, these particular deorbitalized meta-GGAs significantly outperform the explicitly orbital dependent meta-GGA parent (used in gKS form) so far as time per SCF cycle is concerned. While the deorbitalized forms (r2SCAN-L(SRPP) or (SRPP2)) need more cycles to reach self-consistent convergence, the increment is not so large as to offset the much shorter cycle time. As a result, these are much faster than r2SCAN for static lattice solid calculations, as seen in Tables 7, 8 and 9.
@@
That advantage does not propagate into the one AIMD example we have tried. It appears that the density Laplacian dependence in the deorbitalized functionals introduces instabilities that greatly increase the number of SCF cycles needed in the AIMD context. It might seem that the XC potential would be sensitive to small perturbations in the density because of the Laplacian, so that potentials for successive AIMD steps would be different enough to lead to unusual variation in the density, and thus slower SCF convergence. However, our limited testing with reduced AIMD step-time is inconsistent with this argument. Similarly, there is no obvious clue as to the cause of the problematic behavior in the average SCF cycle time versus its spread. Recall that Table 7 shows that r2SCAN deorbitalized by SRPP, SRPP2, and RPP all have similar ratios of spread of time per cycle to average time per cycle, and the ratio for SRPP (2.35) is slightly below that for RPP (2.51).
@@
Because Gaussian basis functions are themselves smooth and are equivalent to fixed combinations of many plane waves, molecules calculated with Gaussian basis sets behave, oddly enough, like the cruder and faster plane-wave calculations, in the sense that there is little distinction in timing among deorbitalizer forms. Utilization of a smooth, analytical basis apparently has a stabilizing effect that is a counterpart to reduction of the energy cutoff in a plane-wave code.
@@
Our limited AIMD calculations also display a third, somewhat discomfiting behavior. As for the case of lower accuracy calculations for the equation of state, deorbitalized functions outperform the orbital-dependent parent on a per-cycle basis, while smoothed and unsmoothed deorbitalizers have similar SCF convergence rates. But PBE and r2SCAN perform the task specific to MD, namely finding orbitals for a new set of nuclear positions starting from the density for a modestly different set of positions, in two or three SCF cycles. In contrast, the deorbitalized forms used here take almost the same number of cycles in starting from the previous converged density for the previous AIMD step as they do when starting from, say, an LCAO density. An attempt to start at each step with a PBE density did not help. The causes are undiagnosed.
@@
Finally, it may be advantageous to use the measure of noise we defined as an aid to construction of deorbitalizers with yet smoother potentials, either by optimizing the switching function between iso-orbital and slowly-varying limits or strategically modifying “appropriate norms" in the deorbitalizer or parent functional. This is the focus of ongoing research.
@@
The following file [40] is available free of charge.
  • SRPP_Timing.SuppInfo: In this file we provide a detailed, system-by-system tabulation of the numerical results of the test calculations against standard molecular and solid test sets. We also provide system-by-system SCF timing and SCF cycle count tabulations, corresponding to the plot of SCF cycle count for each of the first 100 AIMD steps in Fig. 8 above.
@@
Data Availability
The data that support the findings reported in this paper are openly available at the Materials Cloud Archive.[61]